### v2

# BayesSDF: Surface-Based Laplacian Uncertainty Estimation for 3D Geometry with Neural Signed Distance Fields

Abstract
Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps. We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential. Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian splatting, which lack explicit surface formulations, Signed Distance Functions (SDFs) define continuous and differentiable geometry, making them better suited for physical modeling and analysis. BayesSDF leverages a Laplace approximation to quantify local surface instability using Hessian-based metrics, enabling efficient, surfaceaware uncertainty estimation. Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use. Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
1 Introduction
Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations. Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes. Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction. To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field. This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field. In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations. Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function. This formulation ensures that uncertainty maps are geometrically coherent. By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning. The method enables precise identification of areas with high reconstruction variance, leading to improved surface-aware uncertainty estimates. BayesSDF also improves calibration and geometric consistency over previous methods. Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling. This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
2 Contributions
  1. 1.
    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions. This direct integration significantly improves the geometric consistency and calibration of uncertainty predictions.
  2. 2.
    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble or extensive sampling approaches.
  3. 3.
    Scalable and Computationally Efficient Framework: Leveraging a hierarchical hash-based deformation field, BayesSDF maintains high accuracy while reducing computational and memory overhead. The approach integrates seamlessly with existing neural implicit surface frameworks such as NeuS, ensuring a compact yet powerful uncertainty estimation model.
  4. 4.
    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods. Metrics such as the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially in complex real-world scenarios.
  5. 5.
    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical. SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling. Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
These contributions establish BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications for scene reconstruction, physics-based simulations, and autonomous robotic navigation.
3 Related Work
Uncertainty quantification in neural implicit representations is a rapidly evolving field. Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios. In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning. Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, among others. We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
3.1 Uncertainty Estimation in 3D Reconstruction Methods
Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications. Various methods exist for reconstructing 3D surfaces, and each presents unique challenges for uncertainty quantification.
3.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly.
  • NeRF-based methods estimate uncertainty per-ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray. These approaches are effective for rendering applications but fail to capture geometric uncertainty directly.
  • NeuS-based methods instead represent surfaces using a SDF, which provides a continuous and differentiable description of 3D geometry. However, existing NeRF uncertainty estimation techniques do not translate well to NeuS because they rely on ray-space density variations rather than surface-level geometric uncertainty.
3.2 Uncertainty Estimation in Ensemble-Based Learning
Ensemble-based uncertainty estimation has emerged as a powerful technique in deep learning, providing a practical means of approximating epistemic uncertainty. In the context of 3D learning and reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling. While ensemble methods are widely used in 2D vision tasks (e.g. image classification and segmentation), applying them effectively in 3D neural representations presents unique challenges:
  • Computational Cost: Training multiple neural implicit models is memory- and computeintensive, particularly for high-resolution 3D reconstructions. Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
  • Surface Consistency: Unlike 2D image-based tasks, uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays. This results in uncertainty that does not align well with geometric structures, making it less interpretable.
While leveraging the strengths of ensemble-based learning and addressing key limitations in 3D implicit representations, BayesSDF establishes a new paradigm for uncertainty quantification in neural implicit fields.
3.3 Bayesian Inference for Ray-Based Sampling
BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields. It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model. While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty. Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty. Unlike BayesRays, BayesSDF integrates reconstruction methods that inherently reveal volumetric qualities, ensuring direct alignment between uncertainty estimates and the scene geometry. This approach places less emphasis on rendering estimations, making it inherently more interpretable for geometric reconstruction tasks by exploiting surface variability rather than depth-based uncertainty.
Method Representation Uncertainty Estimation Computational Cost Geometric Consistency
NeRF Volume Density None High Low
NeuS SDF None High High
BayesRays Volume Density Ensemble Variance Low Moderate
BayesSDF Volume Density Ensemble Variance Low Very High
Table 1: Comparison of existing 3D reconstruction methods, evaluating their representation, uncertainty estimation, computational cost, and geometric consistency.
4 Background
Quantifying uncertainty in 3D implicit representations is essential for improving the reliability and interpretability of learned geometric models. This section reviews the foundational components upon which BayesSDF builds: Signed Distance Functions (SDFs), Neural Radiance Fields (NeRF), Neural Implicit Surfaces (NeuS), ensemble-based uncertainty methods, and Bayesian estimation frameworks such as BayesRays.
4.1 Signed Distance Functions (SDFs)
Signed Distance Functions (SDFs) define geometry as a continuous field that measures the signed distance of a point in space to the closest surface boundary. Formally, an SDF is defined as $f:\mathbb{R}^{3}\to\mathbb{R}$, where:
$f(\mathbf{x})=\begin{cases}\phantom{-}d,&\text{if }\mathbf{x}\text{ is outsidethe surface}\\ \phantom{-}0,&\text{if }\mathbf{x}\text{ lies on the surface}\\ -d,&\text{if }\mathbf{x}\text{ is inside the surface}\end{cases}$
This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [6]. Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations.
4.2 Neural Radiance Fields (NeRF)
NeRF models learn a volumetric function that maps 3D positions and viewing directions to emitted color and volume density [5]. The function $F_{\theta}:(\mathbf{x},\mathbf{d})\rightarrow(\sigma,\mathbf{c})$ enables photorealistic novel view synthesis by rendering images through volume integration:
[…]
4.3 Neural Implicit Surfaces (NeuS)
NeuS addresses the geometric limitations of NeRF by replacing volumetric density fields with SDFs that model surfaces as zero-level sets [7]. The rendering is adapted by transforming SDF values into opacity via a sigmoid function:
$\alpha(\mathbf{x})=\Phi_{s}(f(\mathbf{x}))=\frac{1}{1+e^{-sf(\mathbf{x})}}$
The final color is computed using transmittance-weighted rendering:
$C(\mathbf{r})=\sum_{i}T_{i}\alpha_{i}\mathbf{c}_{i},\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$
NeuS enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
4.4 Uncertainty Estimation via Ensembles
Ensemble learning estimates epistemic uncertainty by aggregating predictions from multiple independently trained models. For a set of $N$ models, the mean and variance of predictions at point $\mathbf{x}$ are:
$\bar{f}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\mathbf{x}),\quad\text{Var}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}(f_{i}(\mathbf{x})-\bar{f}(\mathbf{x}))^{2}$
This variance captures model disagreement, which is indicative of uncertainty [4]. While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models.
4.5 Bayesian Uncertainty in NeRF: BayesRays
BayesRays proposes a Bayesian approach for NeRF by learning a density-aware variance model and propagating it along rays [2]. The estimated uncertainty is given by:
$U(\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}(\mathbf{x})\cdot T_{k}(\mathbf{x})$
Color and density are modeled as Gaussian random variables, and uncertainty in the rendered image is computed by propagating these distributions through the rendering equation:
$\text{Var}[\hat{C}(\mathbf{r})]=\sum_{i=1}^{N}T_{i}^{2}\left(1-\exp(-\sigma_{i}\delta_{i})\right)^{2}\text{Var}[\mathbf{c}_{i}]$
BayesRays achieves efficient uncertainty estimation without ensembles, but its output lacks surface-level geometric consistency. Integrating such uncertainty with geometric fields remains a core challenge in the field [3].
5 Method
In this section, we detail our proposed BayesSDF method, which leverages a NeuS-like scene representation and adopts an uncertainty estimation framework through a deformation field and a Hessian-based metric. We begin by describing the theoretical underpinnings and motivation, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
5.1 Theory
Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in Red Green Blue (RGB)) rendered by the neural field. Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
5.2 Distortion Fields
The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation. To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding that functions as a voxel-like grid implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, linear interpolation between grid corners, and 3 output channels per queried grid vertex. Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$. This approach is parameter-efficient while remaining flexible. We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid. Specifically, we generate a ray bundle from the dataset’s rays, containing origins, directions, and near/far bounds. We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set). Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray. When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$. Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$. These coefficients are used to blend the deformation vectors from the 8 corners. This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives. Because each point’s deformation influences eight corner entries, we track which corners are involved and in what proportion.
5.3 Calculating the Hessian
The uncertainty computation in BayesSDF revolves around accumulating partial derivatives of the rendered color $\mathbf{c}(\mathbf{x})=(r(\mathbf{x}),\,g(\mathbf{x}),\,b(\mathbf{x}))$ with respect to the deformation parameters stored in the hash grid. Because the deformation field is queried at multiple corners (8 corners per voxel), we track partial derivatives for each corner contribution. Let us denote the final, aggregated color at a particular sample by $\mathbf{C}\;=\;\sum_{\mathbf{x}\in\mathcal{R}\text{-samples}}\mathbf{c}(\mathbf{x})$ where the sum is over all the sample points along the ray (or potentially aggregated over multiple rays). In the code, we backpropagate through each color channel independently to accumulate partial derivatives. For a single color channel $c\in\{r,g,b\}$, we compute $\frac{\partial c}{\partial\mathbf{d}_{c}^{(k)}}$ where $\mathbf{d}_{c}^{(k)}$ refers to the offsets (the deformation parameters) at corner $k\in\{1,\dots,8\}$. We do this by invoking a backwards pass for each color channel and capturing the resulting gradients with respect to $\mathbf{d}$. Each time, we zero out the gradient from the previous backward pass to isolate the new color channel’s derivative. Between each step, the accumulated gradient is saved (and subsequently multiplied by interpolation coefficients, etc.), ensuring that the next backward pass only captures derivatives for the next channel. This pattern also addresses memory constraints: if we did not zero out or separate channel-wise backprop, the gradient buffers for the entire 3D field could become unwieldy. Once these partial derivatives are obtained, we aggregate them in a variable called ‘hessian‘ in the script. Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\Bigl{(}\frac{\partial r}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial g}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial b}{\partial\mathbf{d}_{d}}^{2}\Bigr{)},$
where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ represents each dimension of the deformation vector. In short, we take the partial derivative of r, g, and b with respect to each dimension of the local offset, square them, and sum them to form a scalar representing local “curvature” in color space for that grid corner. We organize these per-corner accumulations into a 3D tensor of size $((2^{\ell})+1)^{3}$. Each cell in that 3D tensor corresponds to a location in the hash-encoded deformation grid. Over multiple training iterations, we repeatedly add these partial-derivative-based values, accumulating them into a global measure that, once normalized or post-processed, can be interpreted as the Hessian with respect to the neural color output and the deformation parameters.
5.4 Distortion Space Perturbations
A critical aspect of BayesSDF is its reliance on small local changes in the deformation field to gauge uncertainty. Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we implicitly examine the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ measure their effect on the rendered color $\mathbf{c}(\mathbf{x})$. In other words, we approximate the local mapping
$\mathbf{c}\bigl{(}\mathbf{x};\,\mathbf{d}(\mathbf{x})+\delta\mathbf{d}(\mathbf{x})\bigr{)}\;\approx\;\mathbf{c}\bigl{(}\mathbf{x};\,\mathbf{d}(\mathbf{x})\bigr{)}\;+\;\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}\bigl{(}\mathbf{x}\bigr{)}\,\delta\mathbf{d}(\mathbf{x}),$
where $\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}(\mathbf{x})$ is precisely the gradient tensor we accumulate in the Hessian array. The magnitude of these partial derivatives quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$. In practice, perturbations are never explicitly applied to $\mathbf{d}(\mathbf{x})$; rather, their effect is simulated by the gradient computation (.backward()). This is central to the BayesSDF mechanism: we do not need to stochastically perturb each voxel corner. Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate local uncertainty.
5.5 Uncertainty Approximation
We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
$\sigma(\mathbf{x})\;=\;\sqrt{\mathbf{H}\bigl{(}\text{grid\_index}(\mathbf{x})\bigr{)}}$
Because $\mathbf{H}$ is computed by adding the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total ’sensitivity’ at a location. One may further process $\sigma(\mathbf{x})$ by normalization or thresholding to highlight regions of highest uncertainty.
6 Results
In this section, we empirically evaluate BayesSDF and demonstrate its effectiveness as a scene reconstruction framework with principled uncertainty estimation. We describe our rendering pipeline, discuss the uncertainty evaluation metrics, and present quantitative and qualitative results, highlighting how BayesSDF is useful for calculating depth accuracy and uncertainty correlation in a geometric sense.
6.1 Rendering Techniques
We employ a custom rendering method that integrates seamlessly with BayesSDF to generate RGB, depth, and uncertainty images under various camera trajectories. This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze uncertainty predictions alongside scene geometry. The method supports producing multiple outputs, most importantly: RGB images, Depth images (visualized with a colormap), and Uncertainty maps (also colormapped, typically “inferno”). Each pixel’s uncertainty value is fed through a user-specified color mapping function to produce interpretive 2D heatmaps of the 3D uncertainty field.
Figure 1: Africa Dataset: BayesSDF-derived view with uncertainty rendering. The left panel shows the RGB image, the middle panel depicts predictive uncertainty, and the right panel illustrates depth.
Figure 2: Basket Dataset: BayesSDF-generated view with uncertainty. The left panel displays the depth map, the middle panel shows the uncertainty estimation, and the right panel presents the RGB image.
The renderer exposes multiple camera path generation strategies:
  • Camera Path Rendering uses externally specified keyframes (e.g., from a JSON file) to produce a smooth trajectory across the scene.
  • Interpolated Rendering creates interpolations between cameras in the training or evaluation datasets, providing a sequence of novel viewpoints.
  • Spiral Rendering generates a spiral sweep around a reference camera pose, which can be useful for capturing 360° animations or for visually inspecting geometry from all sides.
To ensure consistent evaluation, each camera’s rays are generated, forwarded to BayesSDF, and the outputs (rgb, depth, and uncertainty) are concatenated side by side into a single frame. Before rendering, we load a pre-computed Hessian tensor that encodes the local derivatives of color (or equivalently SDF geometry) with respect to the deformation field. BayesSDF’s model is augmented with a custom method that maps these Hessian values to per-pixel uncertainty predictions in the final rendered images. This direct integration allows for real-time overlay of uncertainties without additional overhead.
6.2 Evaluation Metrics
To rigorously validate BayesSDF, we rely on several metrics that capture both reconstruction quality (via error in depth) and the correlation between predicted uncertainty and actual errors. Below are the key metrics, illustrated in our figures and tables.
Figure 3: Africa Dataset: Change in mean absolute error ($\Delta$MAE) as a function of sparsification.
Figure 4: Basket Dataset: Sparsification curve showing the change in mean absolute error ($\Delta$MAE) with increasing levels of sparsification.
Figure 5: Africa Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
Figure 6: Basket Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
  • Pixel Sparsification (x-axis in our plots): We sort the pixels in each rendered view from most uncertain to least uncertain and remove them in cumulative stages. This gradually “filters out” the pixels where the model is presumably the least confident—or equivalently, we can invert the procedure by removing the least uncertain pixels first.
  • Cumulative Mean Absolute Error $\Delta\text{MAE}$ (y-axis): At each sparsification stage, we compute the mean absolute depth error among the remaining pixels in that view. Specifically,
    $\Delta\text{MAE}(p)\;=\;\frac{1}{|\Omega(p)|}\sum_{(u,v)\,\in\,\Omega(p)}\bigl{|}\hat{d}(u,v)-d^{(}u,v)\bigr{|},$
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors. If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ over the entire sparsification curve. A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” indicating the model is effectively ranking pixels by their likelihood of error.
The figures show three colored lines for each dataset/view: the purple line represents sorting by BayesSDF’s uncertainty in descending order, the blue line represents sorting by variance or alternate uncertainty measure (e.g., an ensemble-based variance), and the red line represents the difference between the two curves’ $\Delta\text{MAE}$.
When the red line reaches a low value, it implies that our Hessian-based uncertainty is identifying the same high-error pixels that the variance-based approach finds.
6.3 Analysis
As we navigate the scene via spiral or interpolated trajectories, the uncertainty maps generated by BayesSDF reveal hotspots in regions with fine geometry (e.g., object boundaries, reflective surfaces) or limited camera coverage. Qualitatively:
  • High Uncertainty Regions typically align with complex object contours or occlusion boundaries, matching intuitive expectations.
  • Stable Surfaces (with consistent multi-view coverage) often exhibit low uncertainty, reinforcing that the Hessian-based gradients reflect reconstruction confidence.
[…]
Method Basket Africa Statue Torch
Ensemble 0.206 0.201 0.150 0.164
BayesSDF 0.218 0.268 0.148 0.182
Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) for various datasets and reconstruction methods. Lower AUSE values indicate superior performance.
  • Lower AUSE MAE implies higher correlation between uncertainty and depth error.
  • BayesSDF matches or slightly exceeds ensemble methods on several scenes (notably, “Statue”), while on “Africa” it shows a modestly higher AUSE.
  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training.
Through BayesSDF, we are able to achieve:
  1. 1.
    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local colorspace derivatives with respect to the scene’s underlying deformation field. This approach is both memory and compute efficient.
  2. 2.
    Geometry-Driven Uncertainty: BBy relating uncertainty to the signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric. Regions where small deformation changes cause large color changes correspond to geometry that is poorly constrained or ambiguous.
  3. 3.
    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS rendering pipeline.
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
7 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics. BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements. Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution. BayesSDF enhances these simulations by providing geometric models with quantified uncertainty. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
8 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty. The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality. By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches. Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.
9 Acknowledgments
We would like to thank Shizhan Xu, Wuyang Chen, Daniel Martin, Michael Mahoney, and Jialin Song. We would also like to acknowledge the NERSC and BAIR groups that have backed this project with the necessary resources.